OpenStudy (anonymous):

The probability that an event will happen is 5/8. What is the probability that the event will not happen?

OpenStudy (hockeychick23):

3/8 3/8+5/8=8/8 or 1

OpenStudy (anonymous):

5/8 = 0.625 62.5% that the event will happen 100 - 62.5 = 37.5 37.5% that the event won't happen